OpenStudy (anonymous):

Integrate sin(nx)sin(x) over [0, pi] - Wave Equation

OpenStudy (anonymous):

multiply and divide the given expression by 2 then use the formula 2sin(A)sin(B) = cos(A - B ) - cos (A+B)

OpenStudy (anonymous):

I end up with the answer \[1/2\left[ 1/(n+1)\sin (1+n)x - 1/(n-1)\sin (1-n)x \right]^{\pi}_{0}\] which according to me should be 0 for any value of n

OpenStudy (zarkon):

what if n=1

OpenStudy (zarkon):

\[\int\limits_{0}^{\pi}\sin(x)\sin(x)dx=\frac{\pi}{2}\] so it is not 0 everywhere

OpenStudy (anonymous):

@zarkon for n= 1 it is pi/2 for the rest value of n it is coming out to be zero

OpenStudy (zarkon):

for positive integer n ...I know that